Linear equations using substitution, Thank you !
x/3+y/5 =1
x+6y=12

Answer by jaya_rohith1@yahoo.com(33)

x/3+y/5=1 ----- (1)
x+6y=12 ----- (2)
(1) implies 5x+3y = 15 ------ (3)
(3)*2 implies 10x+6y = 30 ----- (4)
(2) - (4) implies -9x = -18
x = 2
substitute x in (2) we get 2+6y=12
6y=10
y=5/3